OpenStudy (anonymous):

Find the volume of the solid formed by rotating the region enclosed by x=0, x=1, y=0, y=3+x^8 about the x-axis.

OpenStudy (zarkon):

\[\int\limits_{0}^{1}\pi(3+x^8)^2dx\]

OpenStudy (zarkon):

\[=\frac{496\pi}{51}\]